# The Distribution of Planet Radius in Kepler Multiplanet Systems Depends on Gap Complexity

### v1

@@ 5.3 Super-Earth and Sub-Neptune Abundance
The minimum frequency of planet radius in the region of 1-3 R, the bottom of the radius valley, occurs at approximately 1.9 R in our total planet sample. We will define super-Earths as planets larger than Earth and smaller than 1.9 R. Sub-Neptunes will be shorthand for those above 1.9 R and below 3 R. We see the frequency of super-Earths increasing and sub-Neptunes decreasing with increasing gap complexity in Fig. 4. In Fig. 5, super-Earths are more likely to have a high-$C$ while sub-Neptunes are more likely to have a low-$C$.
We calculate the percentage here of each type of planet considering planets only between 1-3 R in our systems divided into $C$$<$0.165 and $C$$>$0.35. In low-$C$ systems, 47% of planets (163 planets) have their median planet radius in the super-Earth regime while 53% (182) are in the sub-Neptune regime. In high-$C$, 63% (103) are super-Earths while only 37% (60) are sub-Neptune. Super-Earths are 1.3 times more prevalent while sub-Neptunes are similarly 1.4 times less prevalent in high-$C$ systems compared to low-$C$ systems.
@@ 6 Discussion
We find that the radius distribution of Kepler planets in 3+ planet systems is correlated with the spacing of planets within a system. While the high-$C$ systems and low-$C$ systems sample the same stellar properties, they sample different spreads in periods. Thus, we discuss if the changes in the radius distribution are a result of period sampling or rather if they could be astrophysical.
@@ 6.1 Sub-Earth Discussion
Furthermore, three of the eight systems have exactly three sub-Earths and three have four or more sub-Earths. Included in these three high-multiplicity sub-Earth systems are KOI-2859 (Kepler-1371) and KOI-4032 (Kepler-1542), two systems of all 5 planets having median planet radius below that of Earth. Additionally, KOI-1151 (Kepler-271) has four out of five candidate planets of less than one Earth-radius. The 6 systems with 3+ sub-Earths that are all in this narrow region of gap complexity (0.07$<$C$<$0.105) have 23 sub-Earths. Removing these 23 planets accounts for about half of the over-abundance of sub-Earths in low-$C$ systems.
While KOI-2859 and KOI-4032 have non-negligible gap complexities, their average $\Delta P/P_{i}$ are 0.29 and 0.26 respectively. In Fig. 1, we see the sub-Earths are primarily under 20 days. To find sub-Earths in multis, the spacing between planets must be small keeping the planets within 20 days. A low spacing is correlated with a low gap complexity as shown in Fig. 3.
This brief analysis of specific systems indicates that the increase in sub-Earths at low-$C$ may be more of an effect of observational bias toward shorter periods for sub-Earth observations than an astrophysical effect. The Kepler telescope can only find sub-Earths with a high number of transits thus favoring shorter orbital periods. With an incomplete survey of small planets, we may be missing systems of sub-Earths with high-$C$.
In addition to missing irregularly spaced systems of sub-Earths, we could also be missing sub-Earths in our already observed high-$C$ systems. Gilbert & Fabrycky (2020) presents several lines of evidence that systems of $C$$>$0.33 have unobserved planets. The value of $C$ for a 4-planet system that is evenly spaced in period ratio where one interior planet is then removed is 0.336. These missing interior planets may not be transiting because of higher mutual inclinations (see next paragraph), but Zhu et al. (2018) find that higher intrinsic multiplicity systems tend to have lower mutual inclinations. Our missing sub-Earths in high-$C$ systems could support that these missing interior planets are small and below the detection threshold. However, the lack of sub-Earths begins at a much lower value of $C$ then suggested in Gilbert & Fabrycky (2020) for missing interior planets with the lowest p-value at $C$$>$0.11.
@@ 6.2 Super-Earth and Sub-Neptune Discussion
Unlike the sub-Earths, we find that change in the abundance of super-Earths and Sub-Neptunes are not an effect of the periods sampled in low- and high-$C$ systems. In Fig. 1, we see there is a small number of hot, near Earth-radii planets which for the most part have high gap complexities. However, if we remove all systems which contain either a shorter than 2 day orbit or longer than 365 day orbit our results remain the same. If this restriction is made there are only 41 systems with $C$$>$0.35, making the uncertainties in the frequency much larger, but the over-abundance of super-Earths and under-abundance of sub-Neptunes remains in high-$C$ systems with approximately the same K-S significance when compared to low-$C$.
The division between low- and high-$C$ does correspond with the value of a system of 3-observed planets and one interior unobserved planet. However, it is less clear why the unobserved interior planet would be preferentially a sub-Neptune over a super-Earth. Nearly two-thirds of high-$C$ systems would need a missing sub-Neptune to equalize the amount of super-Earths and sub-Neptunes. High-$C$ systems do represent a larger spread in periods and sub-Neptunes tend to have longer periods than super-Earths, so there may be a slight bias to missing sub-Neptunes but not enough to match the proportions in low-$C$ systems.
@@
Thus, we too turn our discussion to astrophysical formation mechanisms. If the high-$C$ systems initially had equal amounts of super-Earths and sub-Neptunes, it is possible that the mechanism that raised their gap complexities also changed approximately 10-20% of sub-Neptunes to super-Earths. High-$C$ systems may represent high entropy formation environments which is further supported by He & Weiss (2023) finding cold giant planets in high-$C$ systems. However, Kong et al. (2024) found that the dynamical stirring from giant planets preferentially collide the most different planetesimals in mass and spacing leaving the final system more compact and uniform than in a low-entropy formation environment (also supported by Lammers et al. (2023)).
If planets in high-$C$ systems experience more collisions, this lack of sub-Neptunes could be evidence of impact-driven atmosphere loss. The formation simulations of Dawson et al. (2016) show that systems with low solid surface density and less gas during inner system formation tend to produce rockier planets with wider spacings. Chance et al. (2022) analyzes the simulations of Dawson et al. (2016) for the effect of impact-driven atmosphere erosion and finds it is able to reproduce the radius valley although a less empty radius valley than predicted by photoevaporation. Similarly, the formation models of Izidoro et al. (2022) invoke a period of instability amongst the inner planets in resonant configurations which lead to giant impacts. They argue that the giant impacts are efficient at stripping all primordial atmospheres and the radius valley is between water-rich larger planets and rocky smaller planets.
Alternatively, the difference could form earlier with systems that evolve to high-$C$ initially having less volatile-rich planets. The mechanisms proposed for changing the radius distribution of S-type planets in binary star systems in Sullivan et al. (2023) is the binary decreasing the timescale and mass of the disk which then inhibits core formation. Similarly giant planets, proposed to exist in high-$C$ systems, may decrease the inward pebble drift (Lambrechts & Johansen, 2014) and decrease the mass available in the protoplanetary disk Thommes (2003); Haghighipour & Scott (2012). The decrease of inward pebble drift especially that of volatile-rich material may lead to the overabundance of super-Earths. The shortening of the lifespan of the protoplanetary disk may decrease core masses and leave the planets in more excited orbits which may be unobserved and/or lead to atmosphere erosion by collisions.
@@ 7 Conclusion
The changes in the radius distribution becomes more evident when systems are split into low and high gap complexity samples. A two-sample Kolmogorov-Smirnov (K-S) test confirms that the radii of planets in low-$C$ and high-$C$ systems are drawn from different distributions. The K-S test returns a p-value of 0.0044 for planet radius in systems with $C$$<$0.105 compared to $C$$>$0.11. We find that the difference in the distribution when divided at this value of $C$ is primarily in the sub-Earth regime. Low-$C$ systems below a complexity of 0.105 have 2.5 times more sub-Earths than high-$C$ systems.
When narrowing our analysis to the region of the radius valley between one and three times the radius of Earth, the K-S test returns an even lower p-value (0.0008) when splitting systems into $C$$<$0.165 and $C$$>$0.35. In this case, the planets in low-$C$ systems exhibit a defined radius valley with approximately equal frequencies of super-Earths and sub-Neptunes. High-$C$ systems above a gap complexity of 0.35 have a suppression of sub-Neptune frequency to below that of planets in the radius valley and an increase in the frequency of super-Earths. In high-$C$ systems, 63% of planets have a radius below the radius valley (and above the radius of Earth) and 37% of planets are above the radius valley.
@@
Testing if these formation scenarios can reproduce the radius correlation with gap complexity will be important future work. Significant to understanding these formation scenarios will be future searches for giant planets exterior to inner systems such as in Weiss et al. (2024). A large sample of giant planets exterior to high multiplicity inner systems would allow for the study of how the cold giant planet’s mass, eccentricity, and inclination relate to the properties of the inner system such as gap complexity.
The authors greatly appreciate pointers and feedback from K. Sullivan. This work was inspired by a number of talks and posters at the Extreme Solar Systems V conference in 2024 including those by L. Weiss, K. Sullivan, and B. Liberles. We additionally appreciate M. He, A. Gupta, M. MacDonald, and the Steffen Research Group for stimulating discussions. We acknowledge support from the Astrophysics Research Center at the Open University of Israel, the College of Sciences at the University of Nevada, Las Vegas, and the Nevada Center for Astrophysics. A.V. acknowledges support from the Israel Science Foundation (ISF) grants 770/21 and 773/21.

### v2

@@ 4.3 Super-Earth and Sub-Neptune Abundance
The minimum frequency of planet radius in the region of 1-3 R, the bottom of the radius valley, occurs at approximately 1.9 R in our total planet sample. We will define super-Earths as planets larger than Earth and smaller than 1.9 R. Sub-Neptunes will be shorthand for those above 1.9 R and below 3 R. The planets are colored by categories in Fig. 2. We see the frequency of super-Earths increasing and sub-Neptunes decreasing with increasing gap complexity in Fig. 4. In Fig. 5, super-Earths are more likely to have a high-$C$ while sub-Neptunes are more likely to have a low-$C$.
We calculate the percentage here of each type of planet considering only planets between 1-3 R in our systems divided into $C$$<$0.165 and $C$$>$0.35. In low-$C$ systems, 47% of planets (163 planets) have their median planet radius in the super-Earth regime while 53% (182) are in the sub-Neptune regime. In high-$C$, 63% (103) are super-Earths while only 37% (60) are sub-Neptune. Super-Earths are 1.3 times more prevalent while sub-Neptunes are similarly 1.4 times less prevalent in high-$C$ systems compared to low-$C$ systems.
@@ 5 Discussion
We find that the radius distribution of Kepler planets in 3+ planet systems is correlated with the spacing of planets within a system. While the high-$C$ systems and low-$C$ systems sample the same stellar properties, they sample different spreads in orbital period. Thus, we discuss if the changes in the radius distribution are a result of period sampling or rather if they could be astrophysical.
@@ 5.1 Sub-Earth Discussion
Furthermore, three of the eight systems have exactly three sub-Earths and three have four or more sub-Earths. Included in these three high-multiplicity sub-Earth systems are KOI-2859 (Kepler-1371) and KOI-4032 (Kepler-1542)—two systems of all 5 planets having median planet radius below that of Earth (see Fig. 2). The 6 systems with 3+ sub-Earths that are all in this narrow region of gap complexity (0.07$<$C$<$0.105) have 23 sub-Earths. Removing these 23 planets accounts for about half of the over-abundance of sub-Earths in low-$C$ systems.
While KOI-2859 and KOI-4032 have non-negligible gap complexities, their average $\Delta P/P_{i}$ are 0.29 and 0.26 respectively. In Fig. 1, we see the sub-Earths are primarily under 20 days. To find multiple sub-Earths under 20 days the spacing between planets must be small keeping the planets within 20 days. A tight spacing is correlated with a low gap complexity as shown in Fig. 3.
This brief analysis of specific systems indicates that the increase in sub-Earths at low-$C$ may be an effect of observational bias toward shorter periods for sub-Earth observations. The Kepler telescope can only find sub-Earths with a high number of transits thus favoring shorter orbital periods. The Solar System is an example of a system of sub-Earths with a higher $C$ than the sub-Earth systems found by Kepler. With an incomplete survey of small planets, we may be missing systems of sub-Earths with high-$C$.
In addition to missing irregularly spaced systems of sub-Earths, we could also be missing sub-Earths in our already observed high-$C$ systems. Gilbert & Fabrycky (2020) presents several lines of evidence that systems of $C$$>$0.33 have unobserved planets. The value of $C$ for an intrinsically 4-planet system that is evenly spaced in period ratio where one interior planet is unobserved is 0.336. These missing interior planets may not be transiting because of higher mutual inclinations (see next paragraph), but Zhu et al. (2018) find that higher intrinsic multiplicity systems tend to have lower mutual inclinations. Our missing sub-Earths in high-$C$ systems could support that these missing interior planets are small and below the detection threshold. However, the lack of sub-Earths begins at a much lower value of $C$ than suggested in Gilbert & Fabrycky (2020) for missing interior planets with the deficit most prominent at $C$$>$0.11.
@@ 5.2 Super-Earth and Sub-Neptune Discussion
Unlike the sub-Earths, we find that the change in the abundance of super-Earths and Sub-Neptunes is not easily explained by the periods sampled in low- and high-$C$ systems. In Fig. 1, we see there is a small number of hot, near Earth-radii planets which for the most part have high gap complexities. However, if we remove all systems which contain either a shorter than 2 day orbit or longer than 365 day orbit our results remain the same. If this restriction is made there are only 41 systems with $C$$>$0.35, making the uncertainties in the frequency much larger, but the over-abundance of super-Earths and under-abundance of sub-Neptunes remains in high-$C$ systems with approximately the same K-S significance when compared to low-$C$.
The division between low- and high-$C$ does correspond with the value of a system of 3-observed planets and one interior unobserved planet (discussed in the Sect. 5.1. However, it is less clear why the unobserved interior planet would be preferentially a sub-Neptune over a super-Earth. Nearly two-thirds of high-$C$ systems would need a missing sub-Neptune to equalize the amount of super-Earths and sub-Neptunes. High-$C$ systems do represent a larger spread in periods and sub-Neptunes tend to have longer periods than super-Earths, so there may be a slight bias to missing sub-Neptunes but not enough to match the proportions in low-$C$ systems. An injection-recovery test with varied initial distributions or planet radius and intrinsic gap complexities is reserved for a future work.
@@
Thus, we too turn our discussion to astrophysical formation mechanisms. If the high-$C$ systems initially had equal amounts of super-Earths and sub-Neptunes, it is possible that the mechanism that raised their gap complexities also changed approximately 10-20% of sub-Neptunes to super-Earths. High-$C$ systems may represent high entropy formation environments which is further supported by He & Weiss (2023) finding cold giant planets in high-$C$ systems. In contrast, Kong et al. (2024) find that the dynamical stirring from giant planets preferentially collide the most different planetesimals in mass and spacing leaving the final system more compact and uniform than in a low-entropy formation environment (also supported by Lammers et al. (2023)).
If planets in high-$C$ systems experience more collisions, this lack of sub-Neptunes could be evidence of impact-driven atmosphere loss (Schlichting et al., 2015; Lozovsky et al., 2023). The formation simulations of Dawson et al. (2016) show that systems with low solid surface density and less gas during inner system formation tend to produce rockier planets with wider spacings. Chance et al. (2022) analyzes the simulations of Dawson et al. (2016) for the effect of impact-driven atmosphere erosion and finds it is able to reproduce the radius valley although a less empty radius valley than predicted by photoevaporation. Similarly, the formation models of Izidoro et al. (2022) invoke a period of instability amongst the inner planets in resonant configurations which lead to giant impacts. They argue that the giant impacts are efficient at stripping all primordial atmospheres and the radius valley is between water-rich larger planets and rocky smaller planets.
Alternatively, the difference could form earlier with systems that evolve to high-$C$ initially having less volatile-rich planets. The mechanism proposed for changing the radius distribution of S-type planets in binary star systems in Sullivan et al. (2023) and Sullivan et al. (2024) is the binary decreasing the timescale and mass of the disk which then inhibits core formation. Similarly giant planets, proposed to exist in high-$C$ systems, may decrease the inward pebble drift (Lambrechts & Johansen, 2014) and decrease the mass available in the protoplanetary disk Thommes (2003); Haghighipour & Scott (2012). The decrease of inward pebble drift especially that of volatile-rich material may lead to the overabundance of super-Earths. The shortening of the lifespan of the protoplanetary disk may decrease core masses and leave the planets in more excited orbits which may be unobserved and/or lead to atmosphere erosion by collisions.
@@ 6 Conclusion
Using a sample of 783 transiting planet candidates in 234 systems of three or more planets around FGK stars, we find that the radius of planets in a system correlates with the system’s gap complexity. Specifically, we find:
  • As gap complexity increases, the frequency of sub-Earths ($<$1 R) and sub-Neptunes (1.9$<$R$<$3.0 R) decreases and the frequency of super-Earths (1.0$<$R$<$1.9 R) increases.
@@
  • Systems with $C$$<$0.105 have 2.5 times more observed sub-Earths than systems above $C$$>$0.11.
  • While systems with $C$$<$0.165 have approximately equal numbers of observed super-Earths and sub-Neptunes, systems with $C$$>$0.35 have 1.3 times more super-Earths and 1.4 times fewer sub-Neptunes.
  • A two-sample K-S test indicates that planets in systems with $C$$<$0.105 and in systems with $C$$>$0.11 are not drawn from the same radius distribution below 6 R, with a p-value of 0.004. Similarly, planets in systems with $C$$<$0.165 and in systems with $C$$>$0.35 are not drawn from the same radius distribution between 1-3 R, with a p-value of 0.0008.
@@
Testing if these formation scenarios can reproduce the radius correlation with gap complexity will be important future work. Significant to understanding these formation scenarios will be future searches for giant planets exterior to inner systems such as in Weiss et al. (2024). A large sample of giant planets exterior to high multiplicity inner systems would allow for the study of how the cold giant planet’s mass, eccentricity, and inclination relate to the properties of the inner system such as gap complexity. Lastly, understanding the formation mechanisms that explain our findings is crucial for fitting into the picture the Solar System with its medium gap complexity, rocky planets, cold giant planets, and epoch of giant impacts.
The authors thank the anonymous referee for their feedback which enhanced the quality of this work. We appreciate the pointers and feedback from K. Sullivan. This work was inspired by a number of talks and posters at the Extreme Solar Systems V conference in 2024 including those by L. Weiss, K. Sullivan, and B. Liberles. We additionally appreciate M. He, A. Gupta, M. MacDonald, and the Steffen Research Group for stimulating discussions. We acknowledge support from the Astrophysics Research Center at the Open University of Israel, the College of Sciences at the University of Nevada, Las Vegas, and the Nevada Center for Astrophysics. A.V. acknowledges support from the Israel Science Foundation (ISF) grants 770/21 and 773/21.